# Dieter van Melkebeek

### v1

@@ 3.2 Greedy algorithm
[…]
itemSide effect:
We derive a greedy algorithm for the unit-value setting from the above dynamic program for the general setting. The recursive cases of the Bellman equations (11) can be simplified using the precondition $v\equiv 1$ and be rewritten as
@@
Due to the sorting, Algorithm 4 takes time $\Theta(n\log n)$. For completeness, we point out that the running time can be improved to $\Theta(n)$. The output $S^{*}$ is the longest prefix of the sorted order that has total weight at most $W$. Thanks to the non-negativity of the weights, the length of the prefix $S^{*}$ can be found in linear time using binary search and linear-time selection.
@@ 4 Shortest Paths
Given a digraph $G$, edge lengths, and two vertices, $s$ and $t$, we want to find a shortest path from $s$ to $t$ in $G$, or report that no such path exists. Nonexistence happens if $G$ has no path from $s$ to $t$ at all. When negative edge lengths are allowed, it also happens because there is a path from $s$ to $t$ but every such path can be made shorter. We distinguish between the two cases of nonexistence in the notion of distance.
@@
$d(s,t)\doteq\left\{\begin{array}[]{ll}\infty&\text{ if there is no path from $s$ to $t$}\\ -\infty&\text{ if there is a path from $s$ to $t$ but no shortest one}\\ \min\{\ell(P):P\text{ path }s\rightsquigarrow t\}&\text{ otherwise},\end{array}\right.$
@@
[…]
@@ 4.2 Greedy algorithm
    $d(s,u)+\ell(P)$ for some $u\in\overline{T}$ and some path $P$ inside $\overline{T}$ containing $n-1$ edges. These terms correspond to expansions where no level reaches a vertex $u$ inside $T$.
@@
    The term $d(s,u)+\ell(P)$ of type (c) can be ignored because it corresponds to paths with $P$ as a subpath, and $P$ contains a repeated vertex since it has $n-1$ edges and lies entirely in the set $\overline{T}\subseteq V\setminus\{s\}$ of size at most $n-1$.
@@
we can compute $d(s,v^{*})$ from the distances we already know, namely, as $d(s,v^{*})=d(s,u^{*})+\ell(u^{*},v^{*})$. Thus, we pick $t=v^{*}$ as the next vertex in the order, extend $T$ with $t$ and set $p(v^{*})=u^{*}$.
To efficiently calculate (15), we maintain the vertices $v\in\overline{T}$ in a priority queue, keyed on $\lambda(v)\doteq\min(\{d(s,u)+\ell(u,v):(u,v)\in E\}\cup\{\infty\})$. The iteration for $t$ requires one minimum extraction and $\deg^{+}(t)$ key updates, where $\deg^{+}(t)$ denotes the out-degree of $t$. In total, we need $n\doteq|V|$ minimum extractions and $\sum_{t\in V}\deg^{+}(t)=|E|\doteq m$ key updates. An implementation based on binary heaps runs in time $\Theta((n+m)\log n)$ and an implementation based on Fibonacci heaps in time $\Theta(m+n\log n)$.
@@ 5 Monotonicity Properties
The principle of optimality [Bel54] that underlies dynamic programming can be viewed as a monotonicity property: Better solutions for subproblems cannot worsen the solution to the given instance. The framework of [BdM93] contains an explicit requirement that the combining function of the subproblems be monotone with respect to the objective function; under that condition, the framework produces a correct dynamic program.
@@
[Cur03] also defines local variants of “best” and “better”, where the global objective is replaced by a quality measure for partial solutions. They can be viewed as formalizations of greedy-stays-ahead in the general categorical framework of relational calculus. All four variants enable synthesizing a greedy algorithm within the framework, but need a local criterion $S$ that satisfies the requirement (16) or its alternates. The construction of $S$ and the proof that it works are left to the algorithm designer, as is how to specify a given computational problem so as to allow such $S$.
Our monotonicity properties differ from the ones in the category theoretic setting. In the context of Interval Scheduling and Knapsack, we need to relate the optimal objective values for different subproblems, e.g., for an instance and the instance with one component less. In contrast, the requirements in the category theoretic setting relate the objective values of different solutions to the same subproblem.

### v2

@@ 3.2 Greedy algorithm
$S\subseteq[n]$ such that $|S|$ is maximized and $\sum_{i\in S}w_{i}\leq W$
We derive a greedy algorithm for the unit-value setting from the above dynamic program for the general setting. The recursive cases of the Bellman equations (11) can be simplified using the precondition $v\equiv 1$ and rewritten as
@@
Due to the sorting, Algorithm 4 takes time $\Theta(n\log n)$. For completeness, we point out that the running time can be improved to $\Theta(n)$. The output $S^{*}$ is the longest prefix of the sorted order that has total weight at most $W$. The length of the prefix $S^{*}$ can be found in linear time using binary search and linear-time selection.
@@ 4 Shortest Paths
Given a digraph $G$, edge lengths, and two vertices, $s$ and $t$, we want to find a shortest path from $s$ to $t$ in $G$, or report that no such path exists. Nonexistence happens if $G$ has no path from $s$ to $t$ at all. When negative edge lengths are allowed, nonexistence can also occur because there is a path from $s$ to $t$ but every such path can be made shorter. We distinguish between the two cases of nonexistence in the notion of distance.
@@
[…]
@@
The value version of the problem amounts to computing the distance $d(s,t)$. Typical algorithms implicitly compute the distance for more pairs of vertices than just $(s,t)$. We focus on the single-source variant of the problem, which explicitly outputs the distance for a given source $s$ and all possible targets $t$.
@@ 4.2 Greedy algorithm
    $d(s,u)+\ell(P)$ for some $u\in\overline{T}$ and some path $P$ inside $\overline{T}$ containing n1n-1 edges. These terms correspond to expansions where no level reaches a vertex $u$ inside $T$.
@@
    The term $d(s,u)+\ell(P)$ of type (c) can be ignored because it corresponds to paths with $P$ as a subpath, and $P$ contains a repeated vertex since it has n1n-1 edges and lies entirely in the set $\overline{T}\subseteq V\setminus\{s\}$ of size at most n1n-1.
@@
we can compute $d(s,v^{*})$ from the distances we already know, namely, as $d(s,v^{*})=d(s,u^{*})+\ell(u^{*},v^{*})$. Thus, we pick $t=v^{*}$ as the next vertex in the order, include $t$ into $T$, and set $p(v^{*})=u^{*}$.
To efficiently calculate (15), we maintain the vertices $v\in\overline{T}$ in a priority queue, keyed on $\lambda(v)\doteq\min(\{\infty\}\cup\{d(s,u)+\ell(u,v):(u,v)\in E\})$. The iteration for $t$ requires one minimum extraction and $\deg^{+}(t)$ key updates, where $\deg^{+}(t)$ denotes the out-degree of $t$. In total, we need $n\doteq|V|$ minimum extractions and $\sum_{t\in V}\deg^{+}(t)=|E|\doteq m$ key updates. An implementation based on binary heaps runs in time $\Theta((n+m)\log n)$ and an implementation based on Fibonacci heaps in time $\Theta(m+n\log n)$.
@@ 5 Monotonicity Properties
The principle of optimality [Bel54] that underlies dynamic programming can be viewed as a monotonicity property: Better solutions to subinstances of a given input cannot worsen the solution to the given input. The framework of [BdM93] contains an explicit requirement that the combining function of the subinstances be monotone with respect to the objective function. Under this condition, the framework produces a correct dynamic program.
@@
[Cur03] also defines local variants of “best” and “better”, where the global objective is replaced by a quality measure for partial solutions. They can be viewed as formalizations of greedy-stays-ahead in the general categorical framework of relational calculus. All four variants enable synthesizing a greedy algorithm within the framework, but need a local criterion $S$ that satisfies the requirement (16) or its alternates. The construction of $S$ and the proof that it works are left to the algorithm designer, as is how to (re)formulate a given computational problem in a way that $S$ exists.
Our monotonicity properties differ from the ones in the category theoretic setting. In the context of Interval Scheduling and Knapsack, we need to relate the optimal objective values for different subinstances, e.g., for an instance and the instance with one component less. In contrast, the requirements in the category theoretic setting relate the objective values of different solutions to the same subinstance.